# Capillary waves in miscible fluids

### v1

@@ Abstract
We report the existence of capillary waves propagating along the boundary between miscible co-flowing fluids. We unveil that the interplay between transient interfacial stresses and confinement drives the progressive transition from the well-known inertial regime, characterized by a wavenumber independent of frequency, $k\sim\omega^{0}$, to an unexpected capillary-wave scaling, $k\sim\omega^{2/3}$. This allows us to measure the effective interfacial tension between miscible fluids and its rapid decay on time scales never probed so far, which we rationalize with a model going beyond square gradient theories.
@@
Capillary waves are small disturbances propagating on the interface between two fluids, with such a short wavelength that the only restoring force effectively at work is interfacial tension [1, 2, 3]. They are very common in Nature [1] but they can also be easily produced artificially [4, 5]. Notably, capillary waves have proved to be a valuable tool to measure interfacial or surface tension between immiscible fluids [6], since their dispersion relation links the tension $\Gamma$ to the wavenumber $k$ and frequency $\omega$ of the waves. Recently, Hu and Cubaud [7] have shown that propagating modes can be excited in viscosity-stratified flows confined in microfluidic channels, and that some of the waves observed for coflowing immiscible liquids follow the capillary wave dispersion predicted for inviscid fluids, with a scaling $k\sim\omega^{2/3}$ [1]. Hence, even in confined flows, capillary waves can emerge when interfacial tension dominates over other stabilizing mechanisms such as viscous dissipation. In the flow of miscible fluids interfacial stresses, whose very existence is being debated by an increasing number of research groups in recent years [8, 9, 10, 11, 12, 13], are typically negligible compared to viscous dissipation. An ”inertial” regime characterizes the emerging instability, with $k\sim\omega^{0}$ [7]: the variation of the wave frequency is solely dictated by a change in phase velocity, while the wavelength stays practically unaltered when the flow rate of either fluid is varied. This regime has been also observed for high Reynolds numbers in immiscible fluids, where it delimits the short wavelength limit [7] of the stability diagram, suggesting a maximum wavenumber dictated only by the rate of energy dissipation of the waves. Hence, dissipation, confinement and interfacial tension are key to understand the observed instability and the transition from the inertial to the capillary regime.
@@
Capillary waves in miscible fluids have not been reported so far, most likely due to the rapidly decaying tension at their boundaries [14, 15, 16, 17, truzzolilloOffEquilibriumSurfaceTension2014b, 9]. However, since wave formation at the interface between coflowing liquids in microfluidic devices can be a very fast process, with just a few tens of milliseconds elapsing from the first contact between fluids to the instability occurrence, microfluidics provides an ideal playground to test whether interfacial stresses due to large concentration gradients can transiently impact wave propagation. In this letter, we investigate in detail the propagation of shear flow instabilities appearing when water and glycerol flow side-by-side in rectangular microchannels. We unveil the existence of capillary waves in miscible co-flowing liquids, two distinct branches of wave dispersion, and, from the analysis of wave propagation in space, we measure the effective interfacial tension at the liquid-liquid boundary over time scales never probed so far.
@@
We generated and visualized interfacial waves by pumping deionized water and glycerol in polydimethylsiloxane (PDMS) microchannels, as sketched in Fig. 1 (a) ([19]). Most of the experiments and the detailed analysis of wave propagation have been carried out in a channel whose main duct has height $H=0.1$ mm and width $W=1$ mm. Further experiments to inspect the role of lateral confinement have been performed in a narrower channel, $W=0.25$ mm, with the same height and Y-geometry. The two fluids flow parallel to each other and the interface is vertical, parallel to the gravity direction $\hat{z}$, thus ruling out any effect due to density mismatch, at least in the first portion of the channel, where the two fluids stay in parallel co-flow and do not re-orient themselves to minimize gravitational energy.
@@
The instability is observed via optical microscopy at its onset, and it is characterized by recording the interface dynamics at fixed flow rate of glycerol $Q_{G}$ and varying the flow rate of water from $Q_{H}=0$ $\mu$l/min to several hundreds of $\mu$l/min, depending on $Q_{G}$. We repeated this protocol in the range $3\mu l/min\leq Q_{G}\leq 25\mu l/min$. The instability appears above a critical flow rate of water $Q^{c}_{H2O}$ that depends marginally on $Q_{G}$ ([19]). The typical base flow is sketched in Fig. 1-b, with water being the faster moving fluid. The velocity profiles in absence of diffusion change curvature at the interface, located at $y=Y$. This configuration allows to link the stream wise position $\Delta X$ of the instability onset and the diffusion time $t_{c}$, since $t_{c}=\Delta X/U(Y,H/2)$, where $U(Y,H/2)$ is the base flow velocity of the interface at the imaged plane $z=H/2$.
@@
As shown in Figs 1 (c,d), the interface can be visualized thanks to the large difference in refractive index at T=25 C between water ($n=1.333$) and glycerol ($n=1.472$). Here, both the fluids flow from left to right, with glycerol occupying the upper part of the channel. The shape of the wave front in each frame is reconstructed by tracking the position $Y$ of the interface in the direction orthogonal to the flow direction, $\hat{x}$, thus obtaining a curve in the $x-y$ plane that closely represents the contour of the deformed interface ([19]). The spatio-temporal averages and the respective standard deviations of the wavelength, the phase velocity in the laboratory frame and the amplitude are directly extracted from image analysis ([19]).
@@
Figure 1: Experimental setup and flow configuration. a: Scheme of the Y-junction chip used to generate co-flowing fluids. b: Characteristic 3D-base flow for water and glycerol in the main duct. c: Entire field of view for ($Q_{G}$,$Q_{H}$)=(18 $\mu$l/min, 60 $\mu$l/min) and $W=1$mm. d: Interfacial instabilities observed for three different flow rates of water at $Q_{G}$ = 18 $\mu$l/min. Note that $\lambda(Q_{H}$) is non-monotonic.
@@
Figure 1-(c) and (d) set the scene for describing the general phenomenology of the observed co-flow instabilities. In (c) we show the entire field of view for one single set of flow rates. The instability develops close to the wall at $y=0$. This is always the case in our experiments since the unperturbed interface stays in the range 30 $\mu$m $\leq Y\leq$ 320 $\mu$m. In (d) we show three frames depicting the emerging instability at three selected flow rates of water for the same $Q_{G}$=18 $\mu$l/min. Remarkably, the wavelength of the instability does not vary monotonically upon increasing $Q_{H}$: the average wavelength $\lambda$ of the instability first increases with $Q_{H}$, reaching very large values (up to $\lambda>$W) for intermediate $Q_{H}$, and finally declines when the water rate is increased further. In none of the cases we observe a spatio-temporal evolution of the instability amplitude in the accessible field of view ($1.6$ mm downstream of the onset): almost instantaneously the observed waves reach a steady amplitude, pointing to a non-linear saturation in the instability production with mode-coupling at play [20]. This has been already underlined in [7] for other pairs of fluids.
@@
Figures 2-a,b show the result of our image analysis for the wavelength $\lambda$ and the phase velocity of the waves in the frame of the unperturbed interface, $v_{ph}=v_{ph}^{lab}-U(Y,H/2)$, with $U(Y,H/2)$ computed analytically or retrieved from the interface position ([19]). Note that the subtraction of the base flow velocity at $y=Y$ is necessary to get rid of a Doppler shift caused by the underlying fluid motion [21, 22]. However, we emphasize that our results do not qualitatively change if this subtraction is not considered, as in [7]. As suggested by Fig 1-d the instability wavelength develops a maximum, which grows in magnitude and shifts to higher rates of water for increasing rate of glycerol, Fig. 2-a. By contrast, the phase velocity increases sharply right above the critical rate of water $Q^{c}_{H}$ for the instability (Fig. 2-(b) and [19]), and then becomes very weakly dependent on the water flow rate. This occurs for all the investigated $Q_{G}$. Similarly to the phase velocity, the amplitude of the instability exhibits two regimes with a rapid increase followed by nearly constant plateau as a function of $Q_{H}$ ([19]). Quite generally, $Q_{G}$ determines a systematic increase of the wavelength, the phase velocity and the amplitude ([19]) at any $Q_{H}>Q^{c}_{H}$.
@@
Figure 2: Wavelength $\lambda$ (a) and phase velocity $v_{ph}$ (b) of the observed interfacial waves as a function of the flow rate of water $Q_{H}$, for different flow rates of glycerol as indicated in panel a. The color code and symbol type are the same for both panels. Error bars are spatio-temporal standard deviations ([19]).
Figure 3: Dispersion relation for all $Q_{G}$ from the highest rate (25 $\mu$l/min) to the lowest (3 $\mu$l/min). Blue lines are the fit of the C-branch with a capillary wave scaling $\omega=\alpha k^{3/2}$, see text for details. The C-branch vanishes for $Q_{G}<8$ $\mu$l/min.
@@
Figure 3 shows the dispersion relation $\omega(k)=v_{ph}(k)k$ of the waves, with $k=2\pi/\lambda$. It is characterized by the presence of two distinct branches for all rates of glycerol for which $\lambda(Q_{H})$ is non-monotonic. One can distinguish a first branch, which we call C-branch, corresponding to the highest rates of water, where the frequency increases with increasing $k$, and a second branch, that we call CD, where the frequency decreases with $k$ at high $Q_{G}$, becoming almost $k$-independent and eventually limited over a very small range of $k$ as $Q_{G}$ is decreased down to 3 $\mu$l/min (D-waves), where the C-branch is absent. For the sake of clarity a stability diagram in the $Q_{G}$-$Q_{H}$ plane, including C,CD and D waves, is reported in [19]. Here, C, CD and D stand for capillary, capillary-dissipative and dissipative respectively, as discussed later.
@@
The C-branch is strongly reminiscent of the capillary dispersion observed in immiscible co-flowing liquids [7] and we have thus inspected further this branch, that is indeed well fitted by a capillary wave-type dispersion [1] (blue lines in Fig. 3):
@@
with $\alpha$ that decreases for decreasing $Q_{G}$, until when the C-branch vanishes and only D-waves remain. This scenario is quite surprising since capillary waves in confined shear flows were reported only for immiscible fluids, while miscible ones only showed an inertial regime with a wavenumber independent of frequency [7]. To validate our fitting procedure we report in Fig. 4-a the wavenumber vs the rescaled frequency $\omega/\alpha$ and we fit the entire set of points belonging to the C-branches with a power-law $k=C_{0}(\omega/\alpha)^{\beta}$. We obtain $\beta=0.67\pm 0.05$ in very good agreement with the expected capillary wave scaling $\beta_{c}=2/3$ [1]. From the best fit of eq. 1 to the C-branches we computed the effective tension between water and glycerol for each $Q_{G}$, using the known relationship $\alpha=[\Gamma_{e}/(\rho_{H}+\rho_{G})]^{1/2}$, where $\rho_{H}$ and $\rho_{G}$ are the densities of water and glycerol, respectively [1, 7].
@@
Figure 4-b shows $\Gamma_{e}$ as a function of the average time $t_{c}=\Delta X/\langle v_{int}\rangle$ that the two fluids spend side-by-side until the instability takes place, where $\langle v_{int}\rangle$ is the unperturbed speed of the interface averaged over the rates $Q_{H}$ belonging to the C-branch. Note that this averaging is sound, since the speed of the interface depends mainly on $Q_{G}$ and only very marginally on $Q_{H}$ ([19]). Together with the values extracted from the present analysis of wave propagation, we report also $\Gamma_{e}$ as from the deformation of spinning droplets of water in a reservoir of glycerol 15 s after the first contact between the fluids. The effective interfacial tension shows a very rapid decrease for increasing contact time going from 2.3 mN/m at $t_{c}=$ 61 ms down to 6.3 $\mu$N/m at $t_{c}=$0.17 s, decaying way faster than the $\Gamma_{e}\sim t_{c}^{-1/2}$ scaling predicted for small gradients, where the square gradient (SG) expansion of the free energy is supposed to hold [17, 23]. The values that we obtain are in stark contrast with spinning drop tensiometry data reported previously for the same liquids: in [23] $\Gamma_{e}=0.58$ mN/m after about hundred of seconds from contact, a result leveraging on the assumption that water droplets spinning in a reservoir of glycerol evolve through a series of quasi-equilibrated states during their continuous elongation. The present results question this assumption and call for a different explanation.
@@
We find that our entire set of data, including the one measured in [24], are very well fitted (Fig. 4-b) by a model that some of us developed in [9], where higher order terms in the gradient expansion of the free energy are taken into account. The full expression for $\Gamma_{e}$ reads
$\Gamma_{e}(t_{c})=\Gamma_{SG}+\Gamma_{0}\left[\cosh\left(\frac{L}{\delta(t_{c})}\right)-\frac{L^{2}}{2\delta^{2}(t_{c})}-1\right],$ (2)
where $\Gamma_{SG}$ is the square gradient contribution to the tension, that can be computed on-lattice for molecular fluids [24]:
$\Gamma_{SG}=\frac{RTa^{2}}{V_{m}\delta(t_{c})}\left(\frac{2}{3}+\frac{\chi_{wg}}{6}\right)$ (3)
where $R$ is the ideal gas constant, $a=0.45$ nm and $V_{m}=45.5$ ml/mol are the average diameter and the average molar volume of the fluid molecules, respectively, and $\chi_{wg}=-0.96$ is the Flory interaction parameter for water and glycerol [9]. $\Gamma_{0}=\varepsilon\delta(t_{c})$ sets the scale of the contribution to the effective tension due to large gradients at short times, via the energy density $\varepsilon$ and the interface thickness $\delta(t_{c})$. $L$ is a characteristic length setting the thickness below which higher order terms in the free energy gradient expansion starts affecting the tension. We further assume that the interface thickness $\delta(t_{c})$ increases in time due only to transverse diffusion, $\delta(t_{c})=\sqrt{2D_{wg}t_{c}}$, with $D_{wg}=1.4\times 10^{-11}m^{2}/s$ the diffusion coefficient of water in glycerol [24, 25].
@@
Figure 4: a: Dispersion relation $k$ vs scaled frequency $\omega/\alpha$ for the waves observed in the widest channel ($W=1$ mm). Same color code as in Fig. 1. The shadowed points are those belonging to the CD-branches observed for those $Q_{G}$ where the C-branch exists, and where the amplitude depends on the interface position ([19]). The solid line is the best power-law fit to the C-branch, $k=C_{0}(\omega/\alpha)^{0.67\pm 0.05}$. The dashed horizontal line marks the average value of the wave number, $k_{D}=18mm^{-1}$, obtained by averaging the data for $Q_{G}=$6 $\mu$l and $Q_{G}=$3 $\mu$l. b: Effective interfacial tension $\Gamma_{e}$ vs time elapsed from contact $t_{c}$ (bottom axis) and interface thickness (top axis). The solid line is the best fit obtained via Eq. 2. c: Dispersion relation $k$ vs $\omega$ for the waves observed in the narrowest channel ($W=0.25$ mm). d: Sketch of the general stability diagram proposed for shear flow instabilities at the boundary between 2 liquids in microfluidic flows, with the capillary branch and the inertial regime delimiting the instability region.
Equation 2 therefore contains only 2 free fitting parameters, $L$ and $\varepsilon$, for which we obtain $L=26.5\pm 3.5$ $\mu$m and $\varepsilon=(43\pm 24)\times 10^{-6}$ Pa.
Quite intriguingly, if we assume that the energy density $\epsilon$ is dominated by excluded volume interactions and that local equilibrium [26] may be invoked to use fluctuation-dissipation scaling as crude approximation to estimate the correlation length associated to this density, we obtain $l_{D}=(k_{B}T/\epsilon)^{1/3}=4.54\pm 0.84$ $\mu$m, i.e. a value of the same order of magnitude of the interface thickness in our experiments, marking the transition between the nearly-exponential decay of the tension and the square gradient scaling. Moreover, $l_{D}$ is lower than $L$, in analogy to previous findings for colloidal and polymeric fluids [9]. In this respect it has been argued [27, 28, 29, 30] that the emergence of a characteristic length larger than typical structural length scales is a consequence of a system being strongly driven as observed in a variety of systems, from flowing granular matter, to colloidal and polymer systems. We also note that the steep exponential-like decay of the tension observed for small $t_{c}$ is reminiscent of the steep decay of the tension found in [16] for static conditions and near-critical fluids. Our fit suggests therefore the existence of two distinct decay regimes for interfacial stresses at miscible liquid/liquid boundaries: at short time scales, sharp gradients relax and high-order terms in the free-energy gradient expansion dominate over the square gradient term. At large time scales, by contrast, the concentration gradients are small and the SG term dominates: its contribution is necessary to account for the point that we obtained analyzing spinning drop deformations at large $t_{c}$ (star in Fig 4-b).
@@
We make further use of the best fit (Eq. 2) and from the experimental $t_{c}$ we compute the effective tension for those rates of glycerol for which we do not observe the C-branch. We obtain $\Gamma_{e}=$ 2.97 $\mu$N/m and 1.78 $\mu$N/m for $Q_{Gly}=$6 and 3 $\mu$/min, respectively. We used these values to compute the prefactor $\alpha$ (Eq.1) and report the wavenumber as a function of the rescaled frequency $\omega/\alpha$ also for the two lowest $Q_{G}$ (green symbols in Fig. 4-a). Remarkably, we find that the wavenumber does not show any clear dependence over almost 2 decades in frequency (full range shown in [19]), scattering around an average value $k_{D}=18\pm 2$ mm-1, and merging to the highest wavenumber region of the capillary branch, akin to what has been been observed in [7] for a silicone oil and ethanol, a pair of immiscible fluids with a viscosity contrast similar to glycerol and water (see [19] for a direct comparison).
Our results, in contrast with those reported in [7], thus suggest two different mechanisms driving wave propagation at the boundary between miscible fluids, giving rise to two distinct branches of the wave dispersion. One of them is characterized by the same inertial scaling, the D-branch here, observed in [7], while the other (the C-branch) and the apparent transition between them (the CD regime) are observed in the present work and are absent in [7]. This discrepancy requires a further inspection.

### v3

# Emergence of capillary waves in miscible co-flowing fluids

@@ Abstract
We show that capillary waves can exist at the the boundary between miscible co-flowing fluids. We unveil that the interplay between transient interfacial stresses and confinement drives the progressive transition from the well-known inertial regime, characterized by a frequency independent wavenumber, $k\sim\omega^{0}$, to a capillary wave scaling, $k\sim\omega^{2/3}$, unexpected for miscible fluids. This allows us to measure the effective interfacial tension between miscible fluids and its rapid decay on time scales never probed so far, which we rationalize with a model going beyond square-gradient theories. Our work potentially opens a new avenue to measure transient interfacial tensions at the millisecond scale in a controlled manner.
@@
Capillary waves are small disturbances propagating at the interface between two fluids, with interfacial tension acting as the only relevant restoring force [1, 2, 3]. They are very common in Nature [1] but can also be easily produced artificially [4, 5]. Notably, capillary waves have proved to be a valuable tool to measure interfacial or surface tension between immiscible fluids [6], since their dispersion relation links the tension $\Gamma$ to the wavenumber $k$ and frequency $\omega$ of the waves. Recently, Hu and Cubaud [7] have shown that propagating modes can be excited in viscosity-stratified flows confined in microfluidic channels, and that some of the waves observed for coflowing immiscible liquids follow the capillary wave dispersion predicted for inviscid fluids, $k\sim\omega^{2/3}$ [1]. Hence, even in confined flows, capillary waves can emerge when interfacial tension dominates over other stabilizing mechanisms such as viscous dissipation. In the flow of miscible fluids, by contrast, interfacial stresses are typically negligible compared to viscous dissipation, such that their very existence is debated by an increasing number of research groups [8, 9, 10, 11, 12, 13]. An “inertial” regime characterizes the emerging instability, with $k\sim\omega^{0}$ [7]: the variation of $\omega$ is solely dictated by a change in phase velocity, while the wavelength stays practically unaltered when the flow rate of either fluid is varied. This regime has also been observed for high Reynolds numbers in immiscible fluids, where it delimits the short wavelength limit [7] of the stability diagram, suggesting a maximum wavenumber dictated only by the rate of energy dissipation of the waves. Hence, dissipation, confinement, and interfacial tension are key to understanding flow instabilities and the transition from the inertial to the capillary regime.
@@
To date, the presence of capillary waves in miscible fluids has only been inferred via light scattering in static conditions and for nearly critical fluids [16], whose miscibility can be easily tuned by changing temperature, which however represent a limited subset of fluid pairs to study. The direct observation of capillary waves in fluids, miscible in all proportions at any temperature below their boiling point, has not been reported so far, most likely due to the rapidly decaying tension at their boundaries [14, 15, 16, 17, 18, 9]. Microfluidic devices provides an ideal playground to test whether interfacial stresses can transiently impact wave propagation, since the behavior of the interface between two fluids can be explored on time scales as short as a few tens of milliseconds since the fluids are brought in contact. In this letter, we investigate the propagation of shear flow instabilities appearing when two miscible fluids, water and glycerol, flow side-by-side in rectangular microchannels. We unveil the existence of capillary waves, with two distinct branches of wave dispersion, and we measure the effective interfacial tension at the liquid-liquid boundary over time scales never probed so far, opening a new avenue to measure transient interfacial tensions in a controlled manner at very short times.
@@
We generated and visualized interfacial waves by pumping deionized water and glycerol in polydimethylsiloxane (PDMS) microchannels, as sketched in Fig. 1 (a) ([19]). Most of the experiments have been carried out in a channel whose main duct has height $H=0.1$ mm and width $W=1$ mm. Further experiments to inspect the role of lateral confinement were performed in a narrower channel, $W=0.25$ mm, with the same height and Y-geometry. The two fluids flow parallel to each other and the interface is vertical, parallel to the gravity direction $\hat{z}$, thus ruling out any effect due to density mismatch in the observed region.
@@
The instability is observed via optical microscopy at its onset, and it is characterized by recording the interface dynamics at fixed flow rate of glycerol $Q_{G}$ and varying the flow rate of water from $Q_{H}=0$ $\mu$l/min to several hundreds of $\mu$l/min, depending on $Q_{G}$. We repeated this protocol in the range $3\mu l/min\leq Q_{G}\leq 25\mu l/min$. The instability appears above a critical flow rate of water $Q^{c}_{H2O}$ that depends marginally on $Q_{G}$ ([19]). The typical base flow is sketched in Fig. 1-b, with water being the faster moving fluid. The velocity profiles in absence of diffusion change curvature at the interface, located at $y=Y$. This configuration allows to link the stream-wise position $\Delta X$ of the instability onset and the diffusion time $t_{c}$, since $t_{c}=\Delta X/U(Y,H/2)$, where $U(Y,H/2)$ is the base flow velocity of the interface at the imaged plane $z=H/2$.
@@
As shown in Figs 1 (c,d), the interface can be visualized thanks to the difference in refractive index between water and glycerol. Here, both the fluids flow from left to right, with glycerol occupying the upper part of the channel. The shape of the wave front in each frame is reconstructed by tracking the position $Y$ of the interface in the direction orthogonal to the flow direction, $\hat{x}$, thus obtaining a curve in the $x-y$ plane that closely represents the contour of the deformed interface [19]. The spatio-temporal averages and the respective standard deviations of the wavelength, the phase velocity in the laboratory frame and the amplitude are directly extracted from image analysis [19].
@@
Figure 1: Experimental setup and flow configuration. a: Scheme of the Y-junction chip to generate co-flowing fluids. b: Characteristic 3D-base flow in the main duct, for water and glycerol. c: Entire field of view for ($Q_{G}$,$Q_{H}$)=(18 $\mu$l/min, 60 $\mu$l/min) and $W=1$mm. d: Interfacial instabilities for three different $Q_{H}$, with fixed $Q_{G}$ = 18 $\mu$l/min. Note that $\lambda(Q_{H}$) is non-monotonic.
@@
Figure 1-(c) and (d) set the scene for describing the general phenomenology of the observed co-flow instabilities. In (c) we show the entire field of view for one single set of flow rates. The instability develops close to the wall at $y=0$. Indeed, in all experiments the unperturbed interface stays in the range 30 $\mu$m $\leq Y\leq$ 320 $\mu$m. In (d) we show three frames depicting the emerging instability at three selected flow rates of water for the same $Q_{G}$=18 $\mu$l/min. Remarkably, the wavelength of the instability does not vary monotonically upon increasing $Q_{H}$: the average wavelength $\lambda$ of the instability first increases with $Q_{H}$, reaching very large values (up to $\lambda>W$) for intermediate $Q_{H}$, and finally declines when the water rate is increased further. In none of the cases we observe a spatio-temporal evolution of the instability amplitude in the accessible field of view ($1.6$ mm downstream of the onset): almost instantaneously the observed waves reach a steady amplitude, pointing to a non-linear saturation in the instability production with mode-coupling at play [20]. This has been already underlined in [7] for other pairs of fluids.
@@
Figures 2-a,b show the wavelength $\lambda$ and the phase velocity of the waves in the frame of the unperturbed interface, $v_{ph}=v_{ph}^{lab}-U(Y,H/2)$, with $U(Y,H/2)$ computed analytically or retrieved from the interface position [19]. Subtraction of the base flow velocity at $y=Y$ eliminates the Doppler shift caused by the underlying fluid motion [21, 22]; however, our results do not change qualitatively if no subtraction is performed, as in [7]. As suggested by Fig 1-d, $\lambda$ develops a maximum, which grows in magnitude and shifts to higher rates of water for increasing rate of glycerol, Fig. 2-a. By contrast, the phase velocity initially increases sharply right above the critical rate of water $Q^{c}_{H}$ for the instability (Fig. 2-(b) and [19]), and then nearly plateaus, a behavior seen for all the investigated $Q_{G}$. Similarly to $v_{ph}$, the amplitude of the instability exhibits two regimes with a rapid increase followed by a nearly constant plateau as a function of $Q_{H}$ [19]. Quite generally, the wavelength, the phase velocity and the amplitude [19] increase systematically with $Q_{G}$ at all $Q_{H}>Q^{c}_{H}$.
@@
Figure 2: Wavelength $\lambda$ (a) and phase velocity $v_{ph}$ (b) of the interfacial waves vs the flow rate of water $Q_{H}$, for different flow rates of glycerol as indicated in (a). Same symbols in both panels. Error bars are spatio-temporal standard deviations [19].
Figure 3: Dispersion relation for all $Q_{G}$ from the highest rate (25 $\mu$l/min) to the lowest one (3 $\mu$l/min). Blue lines are fits of Eq. 1 (capillary regime) to the C-branch, which vanishes for $Q_{G}<8$ $\mu$l/min. Error bars are computed propagating the spatio-temporal variances measured for $\lambda$ and $v_{ph}^{lab}$.
@@
Figure 3 shows the dispersion relation $\omega(k)=v_{ph}(k)k$ of the waves, with $k=2\pi/\lambda$. Two distinct branches are seen for all rates of glycerol for which $\lambda(Q_{H})$ is non-monotonic. The first branch, which we term C (capillary), corresponds to the highest $Q_{H}$, with frequency increasing with $k$. In the second branch, termed CD (capillary-dissipative), $\omega$ decreases with $k$ at high $Q_{G}$, becoming almost $k$-independent and eventually limited over a very small range of $k$ as $Q_{G}$ is decreased to 3 $\mu$l/min, where the C-branch is absent and only D (dissipative) waves are seen. See [19] for a stability diagram in the $Q_{G}$-$Q_{H}$ plane, showing the C,CD and D regions. The rationale for the chosen terminology is discussed below.
@@
The C-branch is strongly reminiscent of the capillary dispersion observed in immiscible co-flowing liquids [7]: indeed, this branch is well fitted by a capillary wave-type dispersion [1] (blue lines in Fig. 3):
@@
with $\alpha$ that decreases for decreasing $Q_{G}$, until the C-branch vanishes and only D-waves remain. This scenario is quite surprising since capillary waves in confined shear flows were reported only for immiscible fluids, while miscible ones only showed an inertial regime with a wavenumber independent of frequency [7]. To validate our fitting procedure we report in Fig. 4-a the wavenumber vs the rescaled frequency $\omega/\alpha$ and we fit the entire set of points belonging to the C-branches with a power-law $k=C_{0}(\omega/\alpha)^{\beta}$, obtaining $\beta=0.67\pm 0.05$, in excellent agreement with the capillary wave scaling $\beta_{c}=2/3$ [1]. From the best fit of Eq. 1 to the C-branches we computed the effective tension between water and glycerol for each $Q_{G}$, using the known relationship $\alpha=[\Gamma_{e}/(\rho_{H}+\rho_{G})]^{1/2}$, with $\rho_{H}$, $\rho_{G}$ the densities of water and glycerol, respectively [1, 7].
@@
Figure 4-b shows $\Gamma_{e}$ as a function of the average time $t_{c}=\Delta X/\langle v_{int}\rangle$ that the two fluids spend side-by-side until the instability takes place, where $\langle v_{int}\rangle$ is the unperturbed speed of the interface averaged over the rates $Q_{H}$ belonging to the C-branch. This averaging is sound, since $v_{int}$ depends mainly on $Q_{G}$ and only very marginally on $Q_{H}$ [19]. We also show $\Gamma_{e}$ obtained from the deformation of spinning droplets of water in a reservoir of glycerol [23], 15 s after the first contact between the fluids. The effective interfacial tension shows a very rapid decrease for increasing contact time, from 2.3 mN/m at $t_{c}=$ 61 ms down to 6.3 $\mu$N/m at $t_{c}=$0.17 s, decaying way faster than the $\Gamma_{e}\sim t_{c}^{-1/2}$ scaling predicted for small gradients, where the square gradient (SG) expansion of the free energy is supposed to hold [17, 24], eventually reaching much lower values than those reported previously[24], see [19].
@@
We find that our entire set of data, including the data-point measured in [23], is very well fitted (Fig. 4-b) by a model that some of us developed in [9], where higher order terms in the gradient expansion of the free energy are taken into account. The full expression for $\Gamma_{e}$ reads
$\Gamma_{e}(t_{c})=\Gamma_{SG}(t_{c})+\Gamma_{0}\left[\cosh\left(\frac{L}{\delta(t_{c})}\right)-\frac{L^{2}}{2\delta^{2}(t_{c})}-1\right],$ (2)
[…]
@@
Figure 4: a: $k$ vs scaled frequency $\omega/\alpha$ for the waves observed in the widest channel ($W=1$ mm). Same color code as in Fig. 1. The shadowed points are those belonging to the CD-branches observed for those $Q_{G}$ where the C-branch exists, and where the amplitude depends on the interface position [19]. Error bars are omitted for clarity. The solid line is the best power-law fit to the C-branch detailed in the text. The dashed horizontal line marks the wave number, $k_{D}=18mm^{-1}$, obtained by averaging the data for $Q_{G}=$6 $\mu$l and $Q_{G}=$3 $\mu$l. b: $\Gamma_{e}$ vs $t_{c}$ (bottom axis) and interface thickness (top axis). Black solid line: best fit of Eq. 2. Thick blue line: $\Gamma_{SG}$ computed as detailed in the text and in [19] c: Dispersion relation $k(\omega)$ for the waves observed in the narrowest channel ($W=0.25$ mm). d: Sketch of the general stability diagram.
To fit the entire set of data with Eq. 2, we defined $\delta(t_{c})\equiv\delta_{\nabla}(t_{c})$, the thickness of the region with steepest gradient, delimited by the two points where $\left|\partial^{2}c/\partial z^{2}\right|$ is maximum, and checked that alternative definitions of $\delta$ do not change the scenario described below [19]. We obtained $\delta(t_{c})=\Xi\sqrt{t_{c}}$ with $\Xi=8.24$ $\mu$m/s1/2. (See [19] for the values of $L$ and $\varepsilon$ obtained from the fit and for a comparison with previously reported data). The quantity $t_{c}^{*}=L^{2}$/$\Xi^{2}=$ 19.7 $\pm$ 5.2 s is independent of the definition chosen for $\delta$ and sets a characteristic time scale. For $t_{c}>t_{c}^{*}$, the leading correction to the square gradient expression scales as $t_{c}^{-3/2}$. For $t_{c}<t_{c}^{*}$ higher order terms set in and $\Gamma_{e}(t_{c})$ increases as $\sqrt{t_{c}}\cosh(\sqrt{t_{c}^{*}/t_{c}})$ for $t_{c}\rightarrow 0$.
We find therefore the existence of two distinct decay regimes for interfacial stresses at miscible liquid/liquid boundaries: at short time scales, in the presence of sharp gradients, high-order terms in the free-energy gradient expansion dominate over the square gradient term. At large $t_{c}$, the concentration gradients are small and the SG term dominates, accounting for $\Gamma_{e}$ as obtained by spinning drop tensiometry at large $t_{c}$ (star in Fig 4-b).
@@
We further used the best fit through Eq. 2) and the experimental $t_{c}$ to estimate the effective tension for those rates of glycerol for which no C-branch could be observed, finding $\Gamma_{e}=$ 1.91 $\mu$N/m and 0.96 $\mu$N/m for $Q_{Gly}=6$ and 3 $\mu$/min, respectively. We used these values to compute the prefactor $\alpha$ in Eq.1 and report in Fig. 4 the wavenumber as a function of the rescaled frequency $\omega/\alpha$ also for the two lowest $Q_{G}$ (green symbols). Remarkably, we find that the wavenumber does not show any clear dependence over almost 2 decades in frequency (full range shown in [19]), fluctuating around an average value $k_{D}=18\pm 2$ mm-1, and merging to the highest wavenumber region of the capillary branch, akin to what observed in [7] for silicone oil and ethanol, a pair of immiscible fluids with a viscosity contrast similar to glycerol and water (see [19] for a direct comparison).
Here therefore, the same two branches observed in [7] for a pair of immiscible fluids are also seen for two inter-diffusing liquids, in contrast to the results reported in [7], where the C- and the CD- branches were absent for miscible silicone oils. This discrepancy requires further inspection.